OpenStudy (anonymous):

Input the equation for the following line. Reduce all fractional answers to lowest terms. A line that passes midway between and is parallel to both x - y = 6 and x - y = -6.

OpenStudy (tkhunny):

You may wish to consider the y-intercepts and ponder the meaning of "midway between".

OpenStudy (anonymous):

I am not going to lie, I am still completely lost.

OpenStudy (tkhunny):

No, you're not. If one person is walking down 600 North Street, and another person is walking down 600 South Street, where is the person walking who is exactly between them?

OpenStudy (anonymous):

Obviously zero, but the issue I am having is taking that information and using it to make an equation.

OpenStudy (tkhunny):

No, you just solved the problem. See, it was obvious.

OpenStudy (anonymous):

But 0 is not an equation.

OpenStudy (tkhunny):

It isn't going to bite you. Think on it a bit. x - y = 6 -- y-intercept is what? x - y = -6 -- y-intercept is what? Are these two lines parallel?

OpenStudy (anonymous):

Y intercept is 6 and -6, the lines are parallel.

OpenStudy (tkhunny):

Okay, what do you think of this line? x - y = 0 ??

OpenStudy (anonymous):

Yes, thank you.

OpenStudy (tkhunny):

You can write parallel lines all day. x - y = 3 x - y = 2 x - y = 1 x - y = -1 x - y = 5 x - y = 207 You just have to pick the right one. :-)